OpenStudy (anonymous):

Fan+Medal!! Which equation would best help solve the following problem? The height of a triangle is 6 m more than its base. The area of the triangle is 56 m^2. Find the length of the base.

OpenStudy (cwrw238):

the area of a triangle = half base * height

OpenStudy (anonymous):

h=b+6 area=56^2 area=1/2*b*h

OpenStudy (cwrw238):

area = 56 square meteres if you let length base = x then the height = x + 6

OpenStudy (anonymous):

area =56=\[\frac{ 1 }{ 2 }*base *(base+6)\]

OpenStudy (anonymous):

as height =base+6

OpenStudy (anonymous):

\[56*2=base*(base+6)\]

OpenStudy (anonymous):

is it A?

OpenStudy (anonymous):

112=base^2+6*base base^2+6*base-112=0

OpenStudy (anonymous):

now solve this quadratic equation

OpenStudy (anonymous):

okay I sorta have part two if anyone can help me with that??

OpenStudy (anonymous):

yeah ...Q??

OpenStudy (anonymous):

post the new one

OpenStudy (anonymous):

The height of a triangle is 6 m more than its base. The area of the triangle is 56 m2. Find the length of the base. A. 7 m B. 8 m C. 14 m D. 15 m

OpenStudy (anonymous):

same way it is 56=(1/2)*base*(base+6) base^2+6*base=112

OpenStudy (anonymous):

base^2+6*base-112=0 base^2+14*base-8*base-112=0 base*(base+14)-8*(base+14)=0 (base-8 )*( base+14)=0 base = 8 or base = -14 length can not be negative so base= 8 m

OpenStudy (anonymous):

awesome okay that makes sense!